OpenStudy (anonymous):

evaluate the expression without using a calculator. sin 30 cos 60 + sin 60 cos 30

OpenStudy (anonymous):

sin 30 cos 60 + sin 60 cos 30 Sin 30 = 1/2 cos 60 = 1/2 sin 60 = sqrt(3)/2 cos 30 = sqrt(3)/2 ((1/2)(1/2)) + (sqrt(3)/2)^2 (1/4)+ (sqrt(3)/2)^2

OpenStudy (anonymous):

so would that be the answer?

hartnn (hartnn):

instead u can use sinA cos B +cos A sinB = sin (A+B) so your expression is sin (30+60) = sin 90 = 1

OpenStudy (anonymous):

@franciscanmonk is that my answer?

OpenStudy (anonymous):

@catgirl17 Learn the formula! @franciscanmonk it's great review for you :)

OpenStudy (anonymous):

haha. yeah i thought there was a way to simplify that. and i couldnt get a hold of it.

hartnn (hartnn):

did u get that formula ? @catgirl17 and yes final answer is 1.

OpenStudy (anonymous):

yes